OpenStudy (anonymous):

Will someone please help me with these math questions? When providing the answer, can you also provide an explanation, if possible, so I understand it? Select the coordinates of two points on the line y = -2. (2, -2) and (-2, 2) (-2, -2) and (-2, 0) (2, -2) and (0, -2) (-2, 2) and (-2, -2) Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (cobaltthunder267):

1. Line y=-2 is simply a straight line that intersects the y-axis at -2, therefore, both y intersect points need to be at y=-2. This allows us to eliminate all answers except (c): (2, -2) and (0, -2). 

OpenStudy (anonymous):

Thank you so much, hun! This is really helping me out!

OpenStudy (cobaltthunder267):

2 is a little harder... From converting the given line to slope-intercept form, 4x - 3y = - 12 -3y = -4x - 12 y = - 4/3x - 12 We get the slope, -4/3. Now, we use the equation y-y_1=m(x-x_1) to find the slope-intercept form of the parallel line. Plug in our intercept point (-3, 4) y = m(x - (-3)) Plug in m (the slope): y = -4/3 (x - (-3)) Then add m onto x to get into slope-intercept form: y = -4/3 (x + 3) y = -4/3x + 3 The only problem is that none of your choices has this equation...